There are two cars.First one have V=7.8 m/s.Second one have V=5 m/s.They have same direction but slower one has starter point 38 meters ahead of faster one.Questions:when fast cars will overtake slower?What is the length of the journey they have individually traversed?


1
Expert's answer
2021-10-26T17:18:21-0400
"7.8t=5t+38\\t=13.6\ s\\S=7.8(13.6)=106\ m\\s=106-38=68\ m"